# Mirror: A Multiple-perspective Self-Reflection Method for Knowledge-rich Reasoning

### v1

@@ 6 Conclusion
In this paper, we present a multiple-perspective reflection method, called Mirror, for knowledge-enriched reasoning. To tackle the limitations of LLMs in fact assessment and the generation of high-quality feedback, Mirror is equipped with a directional Navigator, enabling the Reasoner to identify multiple key clues in problem-solving. Furthermore, the consistency among responses generated under different directions enhances the validity of answer assessment, particularly when ground truth is not accessible. Experiments conducted on five reasoning datasets demonstrate Mirror’s superiority over several contemporary CoT-based and self-consistency-based reasoning approaches. Moreover, the ablation study results clearly show that our strategies effectively alleviate the aforementioned challenges.
7 Limitations
@@
8 Ethics Statement
We utilized two publicly available datasets: Massive Multitask Language Understanding (MMLU) and FEVER (Fact Extraction and Verification). MMLU is a multiple-choice question-answering dataset covering 57 subjects across STEM, social sciences, humanities, and more. Notably, some subjects, such as moral disputes and moral scenarios, contain statements raising ethical concerns. LLMs may pose a risk of misuse or misjudgment in these contexts. We strongly recommend thorough consideration of safety implications before applying relevant techniques in real-world scenarios. For the FEVER dataset, positive claims (facts) are extracted from Wikipedia, and negative claims are generated by contrasting these facts and subsequently verified without knowledge of their source sentences. However, considering Wikipedia as a social network where virtually anyone can edit content, the extracted facts may not always be flawless. Consequently, we advise against utilizing our work as the sole ground truth for any fact verification task to avoid potential confusion and bias.
@@ Appendix B Mirror algorithm
The overall process consists of three steps: (1) SearchPolicy to obtain the terminal node $v_{l}$. through which expands the tree until fully expanded. Specially, we randomly add one or more nodes to the root node according to the possible actions. In our case, we generate multiple responses to the given question and previous attempts/response. When the current node is fully expanded, we apply the UTC algorithm to select the best child node. (2) Simulation the reward $r$ for $v_{l}$ through SimulationPolicy. This phrase is to simulate the future rewards of the current node through multiple interactions. For simplicity, we follow the similar process as expansion and return the reward $r$ for selected action-state pair. (3) BackPropagate the simulation results to the selected nodes to accelerate SearchPolicy in next iteration.
@@
state transition function $f\mathrel{\mathop{:}}S\times A\rightarrow S$, weight $C_{p}$, Reward $\mathcal{R}$, Stop Criteria $g\rightarrow\{0,1\}$
@@ C.1 Implementation Details
In order to encourage diverse direction generation, we set the generation temperature as 0.8 for all the models, and we set $\texttt{do}_{\_}\texttt{Sample}=\texttt{True}$ for llama and vicuna to avoid greedy search. For the threshold $T_{0}$ in self-assessment to deriving the final answer, we set 0.8 for GPT35, and 0.5 for llama and Vicuna according to the results on limited validation data.
@@
Prompt for Direction Generation (MMLU)
Prompt for Direction Generation (FEVER)
Prompt for Response Generation (MMLU)
Prompt for Response Generation (FEVER)

### v2

@@ 6 Conclusion
In this paper, we present a multiple-perspective reflection method, called Mirror, for knowledge-enriched reasoning. To tackle the limitations of LLMs in fact assessment and the generation of high-quality feedback, Mirror is equipped with a directional Navigator, enabling the Reasoner to identify multiple key clues in problem-solving. Furthermore, the consistency among responses generated under different directions enhances the validity of answer assessment, particularly when ground truth is not accessible. Experiments conducted demonstrate Mirror’s superiority over several contemporary CoT-based and self-consistency-based reasoning approaches without access to ground truth. Moreover, the ablation study results clearly show that our strategies effectively alleviate the aforementioned challenges.
Limitations
@@ Appendix B Mirror algorithm
The overall process consists of three steps: (1) SearchPolicy to obtain the terminal node $v_{l}$. through which expands the tree until fully expanded. Specially, we randomly add one or more nodes to the root node according to the possible actions. In our case, we generate multiple responses to the given question and previous attempts/response. When the current node is fully expanded, we apply the UTC algorithm to select the best child node. (2) Simulation the reward $r$ for $v_{l}$ through SimulationPolicy. This phrase is to simulate the future rewards of the current node through multiple interactions. For simplicity, we follow the similar process as expansion and return the reward $r$ for selected action-state pair. (3) BackPropagate the simulation results to the selected nodes to accelerate SearchPolicy in next iteration. Algorithm is described in Algorithm 1.
@@
state transition function $f:S\times A\rightarrow S$, weight $C_{p}$, Reward $\mathcal{R}$, Stop Criteria $g\rightarrow\{0,1\}$
@@ C.1 Implementation Details
In order to encourage diverse direction generation, we set the generation temperature as 0.8 for all the models, and we set $\texttt{do}_{\_}\texttt{Sample}=\texttt{True}$ for llama and vicuna to avoid greedy search. For the threshold $T_{0}$ in self-assessment to deriving the final answer, we set 0.8 for GPT35, and 0.5 for llama and Vicuna according to the results on limited validation data. These results reveal that larger language models are more consistent in their multiple outputs, which is more difficult for the smaller models. Hence, we adopt a relatively lower threshold for smaller models. This observation can be partially explained by the tendency of larger LMs to rely on their parametric memory (Xie et al., 2023a).
@@
Prompt for Direction Generation (MMLU) As a tutor, your focus is on guiding the student to navigate multiple-choice question-answering problems strategically. Encourage them to dissect the question, identifying key elements and nuances within each choice. Emphasize the importance of understanding subtle differences that could distinguish correct from incorrect options. As a tutor, your are supposed to meticulously evaluate the student’s approach to multiple-choice problems. Question, Choices and the student’s previous thought and answer are given, check if the facts mentioned in the thought is correct and if there might be a more appropriate option than the one chosen. If the student’s reasoning thought is accurate and the proposed answer is the most appropriate, encourage them to adhere to their initial trial. Otherwise, guide the student to revisit specific details, explore alternative choice.
Prompt for Direction Generation (FEVER) As a tutor, your focus is on guiding the student to navigate fact-checking problems strategically. Encourage them to dissect the claim, identifying key elements and associate facts. Emphasize the correct relation between important elements that could distinguish SUPPORTS from REFUTES options. Also, lacking of enough information will lead to NOT ENOUGH INFO. As a tutor, your are supposed to meticulously evaluate the student’s approach to fact verification task. Claim and the student’s previous thought and answer are given, check if the relations mentioned in the Thought is correct and if there might be a more appropriate answer. If the student’s reasoning thought is accurate and the proposed answer is the most appropriate, encourage them to adhere to their initial trial. Otherwise, guide the student to revisit specific details, explore alternative answer.
Prompt for Response Generation (MMLU) You are an expert in multiple-choice question answering. Each problem will provide you with a question and answer choices. Read the question and all the choices carefully, along with the provided advice, and solve the problem by having a thought. Thought can reason about the current situation. Finish[answer] returns the answer and finishes the task. You’re an advanced reasoning agent capable of self-reflection and continuous improvement. Your objective is to tackle multiple-choice question answering problems. Each problem will provide you with a question, answer choices, your previous line of reasoning, and the detailed analyses from an experienced tutor. In a succinct review, assess the accuracy of your earlier answer based on your expertise and the advice, subsequently arrive at the definitive response.
Prompt for Response Generation (FEVER) You are a knowledgeable and accurate fact verifier. Please verify the correctness of the following claim based on your expertise and provided advice. Return SUPPORTS or REFUTES a Claim, or if there is NOT ENOUGH INFO. You’re an advanced reasoning agent capable of self-reflection in fact verification task. Claim and the your previous response and answer are given, along with the advice. In a succinct review, assess the accuracy of your earlier answer based on your expertise and the advice, subsequently arrive at the definitive response.
@@ C.2 Additional Results
We provide additional results as complementary to our main results.
Results on GSM8K dataset.
To further verify the effectiveness of Mirror in various reasoning tasks, we include the math problem, i.e., GSM8k (Cobbe et al., 2021). The performance superiority is evident when comparing with the best-performing unsupervised baseline, i.e., self-consistency .
[…]
0.72 0.75 0.77 0.80
Table A4: Based on GPT-3.5, performances of different reasoning methods.
@@
To analyse the effects of different LLMs quantitatively, we calculate the average pairwise semantic similarity between multiple directions for one question, then 1- similarity to obtain the diversity measurement shown below. The pretrained model, all-MiniLM-L6-v2  [^7] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 is used for sentence-pair similarity calculation. The results is consistent with the intuition that sophisticated LLMs incline to less diverse instruction although such diverse directions are already capable to improve task performances.
Stem Social Humanity Other Fever
GPT-3.5 0.665 0.673 0.674 0.648 0.632
LLama 0.689 0.691 0.688 0.669 0.661
Vicuna 0.694 0.690 0.692 0.679 0.673
Table A5: Diversity among different directions generated by Mirror.
Appendix D Ethics Statement
We utilized two publicly available datasets: Massive Multitask Language Understanding (MMLU) and FEVER (Fact Extraction and Verification). MMLU is a multiple-choice question-answering dataset covering 57 subjects across STEM, social sciences, humanities, and more. Notably, some subjects, such as moral disputes and moral scenarios, contain statements that may raise ethical concerns. Here, LLMs could be misused or misinterpret the information. We strongly recommend thorough consideration of safety implications before applying such techniques in real-world scenarios. For the FEVER dataset, positive claims (facts) are extracted from Wikipedia, and negative claims are generated by contrasting these facts and subsequently verified without knowledge of their original source sentences. However, due to Wikipedia’s editable nature, the extracted facts may not always be entirely accurate. Consequently, we advise against solely relying on our work as the truth source for any fact-checking task to avoid potential confusion and bias.